OpenStudy (anonymous):

use general power rule to find the derivative of \[(2x-3)x ^{3/2}\]

OpenStudy (anonymous):

are you taking calc too, i think i might know how to do this one,what chapter is it?

OpenStudy (anonymous):

Yes I am. Still having trouble with derivatives though. If you know how to do it I'd love to know. In my book it's chapter 3 section 8

OpenStudy (anonymous):

It Can be done by using Product Rule. And the answer is \[2x^{\frac{ 3 }{2}}+(2x-3)\frac{ 3 }{2}\sqrt{x}\]

OpenStudy (anonymous):

yea that's what i got but does that simplify further into \[\frac{ \sqrt{x} }{ 2 }(10x-9)\] because that's the answer in the back of my book